Machine X has an initial cost of $14,000. It is expected to last 10 years, to cost $255 per year to maintain and to have a salvage value of $6,000 at the end of its useful life. The equivalent uniform annual cost of the machine at 7% interest is approximately _____________.
Answer #1

. The equivalent uniform annual cost of the machine at 7% interest is

EUAC = 14,000(A/P, 7%, 10) + 255 - 6,000(A/F, 7%, 10) [ from table, we get values]

= 14000*0.1424 + 255 - 6000*.0724 = 1,814.2